OpenStudy (pokemon23):

If y = 5x + 2, then find the value of 10xy + 4y in terms of y.

OpenStudy (pokemon23):

@Hero @Calcmathlete

OpenStudy (bahrom7893):

wow u didn't type @bah here...

OpenStudy (bahrom7893):

I'm not smart enough????

OpenStudy (bahrom7893):

y=5x+2 y-2 = 5x x = (1/5)(y-2)

OpenStudy (bahrom7893):

10xy + 4y = 10 (1/5)(y-2) y + 4y = 2(y - 2)y + 4y = 2y^2 - 4y + 4y = 2y^2 <-Final answer

OpenStudy (bahrom7893):

and it's you're

OpenStudy (pokemon23):

o_o I'm taking the PSAT... I don't know how to do these types of questions......

OpenStudy (anonymous):

To clarify on what @bahrom7893 said, I would've done something a bit simpler? \[\implies10xy + 4y~~~~~~~~~~Original~Expression\]\[\implies 2y(5x + 2)~~~~~~~~~Factor~out~the~GCF\]\[\implies 2y(y) ~~~~~~~~~~~~~~~~~~Substitute\]\[\implies 2y^2~~~~~~~~~~~~~~~~~~~~~Multiplication~of~variables\]

OpenStudy (pokemon23):

ok let's do another problem I need to get better.

OpenStudy (anonymous):

Uh ok then... \[y = (3x + 2)^3\]\[\text{Simplify in terms of y: } (9x + 6)^3\]

OpenStudy (anonymous):

Very similar to the one you were doing on that PSAT question.

OpenStudy (pokemon23):

ok I think I got it.

OpenStudy (pokemon23):

so do I multiply to the 3rd power?

OpenStudy (anonymous):

Yes, the \(^{'3'}\) implies cubing it or multiplying by itself 3 times.

OpenStudy (pokemon23):

9x*9x*9x?

OpenStudy (anonymous):

Not quite...could you show me what you did?

OpenStudy (pokemon23):

well you its  so i multiply the exponent to the 3rd?